OpenStudy (anonymous):

Help! PLEASE! The length of a rectangle is 3ft more than twice its width, and the area of the rectangle is 54ft^2. Find the dimensions of the rectangle

OpenStudy (dean.shyy):

The rectangles perimeter should be w + w + (l + 3) + (l+3). So, the area should be w * (l+3) = 54 feet squared. Now calculate the dimensions using this information.

OpenStudy (anonymous):

Call the width 'x'. Width = x Length = x + 3 Area = Length x Width = x(x+3) = x^2 + 3x = 54 Since we know x^2 + 3x = 54, bring 54 over and solve the quadratic to get values of x. x^2 + 3x - 54 = 0 (x + 9)(x - 6) = 0 x = -9, 6 Since we know that x represents the width of the rectangle, it must be a positive so we scrap the negative solution. Therefore, our solution for the width is x = 6. We know that the length = x + 3 = 6 + 3 = 9. Has, the dimensions are length = 9 width = 6, 9 by 6. @maddiemaze